Donald, Anne, and Todd have the following capital balances; $40,000, $50,000 and $30,000 respectively. The partners share profits and losses 20%, 40%, and 40% respectively. What is the total partnership capital after Anne retires receiving $80,000 and using the bonus method?(A:40,000)
Answer #1

total partnership capital after. Anne retires receiving \(\$ 80,000\) and using the bonus method

Ans- 40000

The \(\$ 30,000\) bonus is deducted from the remeining partners accordingto their relative profit and loss ratio. Donald \(=20 \%\) and Todd \(=40 \%\) which is a \(1 / 3,2 / 3\) split, Donald \(=\$ 40,000\) \((1 / 3 \times \$ 30,000)=\$ 30,000 .\) Todd \(=\$ 30,000-(2 / 3 \times \$ 30,000)=\)

\(\$ 10,00 d\)